# FP4DiT: Towards Effective Floating Point Quantization for Diffusion Transformers

### v1

@@ 5 Conclusion
Supplementary Material
A Proof for Scale-Aware AdaRound
AdaRound [34] and BRECQ [24] use gradient descent to update the rounding mask. According to Eqs. 4, 5 and 6, ignoring the regularizer $f_{\mathrm{reg}}(\mathbf{V})$, we have following theorem which provides theoretical evidence that the origin AdaRound is imbalance across different scales $s$.
@@ A Proof for Scale-Aware AdaRound
The learning objective doesn’t change for scale-aware AdaRound. Hence, from Eq. 4 and 7, we give:
@@
A.1 Additional Ablation Studies
@@
A.1.1 Effect of FP Format.
A.1.2 Effect of Group Size.
We compare the group-wise weight quantization result with different group sizes in Table 7. Decreasing the group size $g$ for weight quantization consistently improves the quantization performance down to $g=128$. According to [40], group size $g$ such as 128 can result in substantial improvement while maintaining low latency.
@@
0:   Full-precision array $A_{\text{FP}}$, number of bits $n$, number of exponent bits $n_{e}$, number of mantissa bits $n_{m}$, clipping value $maxval$
   $A_{\text{abs}}\leftarrow\text{abs}(A_{\text{FP}})$
   $bias\leftarrow 2^{n_{e}}-\text{log}_{2}(A_{\text{abs}})+\text{log}_{2}(2-2^{-n_{m}})-1$
   $A_{\text{clip}}\leftarrow\text{min}(\text{max}(A_{\text{FP}},-maxval),maxval)$
   $S_{\text{log}}\leftarrow\text{clamp}(\lfloor\text{log}_{2}(\text{abs}(A_{\text{clip}}))\rfloor+bias)),1)$
   $S\leftarrow 2.0^{(S_{\text{log}}-n_{m}-bias)}$
   $result\leftarrow\text{round-to-nearest}(A_{\text{clip}}/S)\times S$
   return  $result$
B Extended Experimental Settings
B.1 Floating Point Quantization Scheme
Since Floating Point Quantization (FPQ) is not as straightforward as INT quantization, there hasn’t been a simple and unified algorithm for performing FPQ yet. In this paper, we apply Algorithm 1 from [20] to perform the FPQ. Unlike [20] learning the clipping value $maxval$ and bit allocations between mantissa and exponent part, we use the absolute maximum of the tensor as $maxval$ and perform FPQ with a predetermined FP format.
@@ B.1 Floating Point Quantization Scheme
B.2 Baseline implementation
In this paper, we compare FP4DiT with three baseline: Q-Diffusion [23], TFMQ-DM [15] and ViDiT-Q [58]. Their implementation details are enumerated as follows:
  1. 1.
    Q-Diffusion: Q-Diffusion is developed on DDIM, LDM, and the original Stable Diffusion v1.4 repository but does not provide native support for any DiT-based DM. Thus, we integrate the Hugging Face Diffusers [50] API with Q-Diffusion to provide support for PixArt-$\alpha$ and PixArt-$\Sigma$. Q-Diffusion has two techniques tailored for Diffusion Models (DM): time step-aware calibration and shortcut-splitting quantization. However, the shortcut-splitting is unworkable for the PixArt model as there are no U-Net shortcuts in either PixArt DiT. As for the activation quantization, we replaced their temporal-aware calibration quantization with online token-wise quantization for better performance and fair comparison.
  2. 2.
    TFMQ-DM: TFMQ-DM is also originally developed for U-Net DMs so we integrate the Hugging Face Diffusers API. However, its Temporal Information Block (TIB) consolidates all embedding layers and time embed into a unified quantization block, which is based on the unique structure of U-Net. Our solution is to replace the TIB in TFMQ-DM with the time embedding module in the DiT models. Similar to Q-Diffusion, online activation is also applied to the TFMQ-DM.
  3. 3.
    ViDiT-Q: ViDiT-Q supports the quantization PixArt-$\alpha$ and PixArt-$\Sigma$. ViDiT-Q uses a mix-precision quantization strategy: For certain layers, ViDiT-Q quantizes them with higher precision like 8 bits, or skips their quantization (16 bits). In our experiments, we use it as W4 and perform the comparison. Besides, they employ online activation quantization, which can be fairly compared to FP4DiT.
Benchmark HPSv2 MS-COCO
[…]
PixArt-$\alpha$ Full Precision W16A16 32.56 31.06 30.76 29.67 31.01 50.42 0.3075
Q-Diffusion W4A8 24.18 23.43 22.91 22.15 23.17 70.34 0.3017
TFMQ-DM W4A8 27.88 26.03 25.14 24.91 25.99 72.41 0.3066
ViDiT-Q W4A8 17.93 17.35 16.81 17.59 17.42 49.04 0.2900
FP4DiT (ours) W4A8 28.63 26.79 27.59 26.72 27.43 46.18 0.3076
Q-Diffusion W4A6 12.63 13.39 13.32 10.65 12.50 92.24 0.2868
TFMQ-DM W4A6 24.02 23.36 22.72 23.04 23.29 75.24 0.3015
ViDiT-Q W4A6 16.71 16.28 16.23 16.56 16.44 57.81 0.2827
FP4DiT (ours) W4A6 24.57 23.08 23.30 23.26 23.55 57.17 0.3031
PixArt-$\Sigma$ Full Precision W16A16 33.07 31.58 31.54 30.49 31.67 47.97 0.3139
Q-Diffusion W4A8 27.30 26.11 26.06 25.06 26.13 50.63 0.3050
TFMQ-DM W4A8 24.95 23.59 23.19 22.73 23.62 77.82 0.2975
ViDiT-Q W4A8 27.10 26.45 26.24 24.49 26.07 50.43 0.2562
FP4DiT (ours) W4A8 27.95 26.29 26.16 24.67 26.27 48.73 0.3064
Q-Diffusion W4A6 24.07 22.42 22.47 21.72 22.67 60.25 0.3027
TFMQ-DM W4A6 19.30 18.46 18.85 17.50 18.53 169.24 0.2346
ViDiT-Q W4A6 24.84 23.23 23.48 21.94 23.37 106.22 0.2425
FP4DiT (ours) W4A6 26.91 25.55 25.22 23.93 25.40 58.59 0.3040
Hunyuan Full Precision W16A16 33.72 31.84 31.52 31.24 32.08 68.41 0.3102
Q-Diffusion W4A8 26.00 24.74 24.84 24.26 24.96 90.16 0.3006
TFMQ-DM W4A8 28.77 27.63 27.67 26.15 27.56 84.34 0.3075
FP4DiT (ours) W4A8 28.96 27.75 27.47 26.94 27.78 79.51 0.3102
Q-Diffusion W4A6 14.90 14.83 15.32 14.24 14.83 264.75 0.2277
TFMQ-DM W4A6 13.73 13.31 13.16 14.51 13.68 256.42 0.2520
FP4DiT (ours) W4A6 15.23 14.94 15.57 14.27 15.00 248.11 0.2562
Table 9: Quantitative evaluation results for PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan in terms of FID and CLIP score. Specifically, for each configuration, we generate 10k images (3k for Hunyuan) using COCO 2017 validation set prompts. Best and second best results in bold and italics, respectively.
B.3 Evaluation Settings
For CLIP score [12], we apply the openai-clip  [^1] https://github.com/openai/CLIP library to measure the CLIP score between prompts and generated images. We employ ViT-B/32 as the CLIP model. To measure HPSv2 [52] score, we generate 3.2k images across the four categories (800 images per category) and use the provided human preference predictor to estimate the performance. For the Fréchet Inception Distance (FID) [13] measurement, we apply clean-fid  [^2] https://github.com/GaParmar/clean-fid library to measure the FID between generated images and ground-truth images.
B.4 Hardware and Software Resources
@@
Our code is running under Python 3 using Anaconda virtual environments and open-source repository forks based on Q-Diffusion [23]. We modified the code to implement our FP4DiT and enable the interface between Quantization scripts and the Hugging Face Diffusers  [^3] https://huggingface.co/docs/diffusers/en/index library. The hardware cost measurement is conducted using the ptflops  [^4] https://github.com/sovrasov/flops-counter.pytorch library. We provide a code implementation with README listing all necessary details and steps.
C Additional Quantization Results
In this section, we present the detailed HPSv2 scores and the FID and CLIP on the MS-COCO dataset. Table 9 enumerates all quantitative results for FP4DiT and other baselines on PixArt-$\alpha$, PixArt-$\Sigma$ and Hunyuan. Aligning with FP4DiT’s superior HPSv2 score shown in Table 2, FP4DiT outperforms all baseline methods across different precisions and models in terms of FID and CLIP. Although FP4DiT doesn’t lead under some HPSv2 categories, FP4DiT’s score is close to the best thus achieve best average HPSv2 score across all tasks. The consistent superior results on HPSv2, FID and CLIP verify the effectiveness of FP4DiT.
D Additional Visualization Results

### v2

@@ 5 Conclusion
Appendix A Appendix
A.1 Proof for Scale-Aware AdaRound
AdaRound (Nagel et al., 2020) and BRECQ (Li et al., 2021) use gradient descent to update the rounding mask. According to Eqs. 4, 5 and 6, ignoring the regularizer $f_{\mathrm{reg}}(\mathbf{V})$, we have following theorem which provides theoretical evidence that the origin AdaRound is imbalance across different scales $s$.
@@ A.1 Proof for Scale-Aware AdaRound
The learning objective does not change for scale-aware AdaRound. Hence, from Eq. 4 and 7, we give:
@@
A.2 Additional Ablation Studies
@@
A.2.1 Effect of Group Size.
We compare the group-wise weight quantization result with different group sizes in Table 9. Decreasing the group size $g$ for weight quantization consistently improves the quantization performance down to $g=128$. According to Park et al. (2022), group size $g$ such as 128 can result in substantial improvement while maintaining low latency.
@@
1:Full-precision array $A_{\text{FP}}$, number of bits $n$, exponent bits $n_{e}$, mantissa bits $n_{m}$, clipping value $maxval$
2:$A_{\text{abs}}\leftarrow\text{abs}(A_{\text{FP}})$
3:$bias\leftarrow 2^{n_{e}}-\log_{2}(A_{\text{abs}})+\log_{2}(2-2^{-n_{m}})-1$
4:$A_{\text{clip}}\leftarrow\min(\max(A_{\text{FP}},-maxval),maxval)$
5:$S_{\text{log}}\leftarrow\text{clamp}(\lfloor\log_{2}(\text{abs}(A_{\text{clip}}))\rfloor+bias,1)$
6:$S\leftarrow 2.0^{(S_{\text{log}}-n_{m}-bias)}$
7:$result\leftarrow\text{round-to-nearest}(A_{\text{clip}}/S)\times S$
8:return $result$
A.3 Extended Experimental Settings
A.3.1 Floating Point Quantization Scheme
Since Floating Point Quantization (FPQ) is not as straightforward as INT quantization, there has not been a simple and unified algorithm for performing FPQ yet. In this paper, we apply Algorithm 1 from (Kuzmin et al., 2022) to perform the FPQ. Unlike (Kuzmin et al., 2022) learning the clipping value $maxval$ and bit allocations between mantissa and exponent part, we use the absolute maximum of the tensor as $maxval$ and perform FPQ with a predetermined FP format.
@@ A.3.1 Floating Point Quantization Scheme
A.4 Evaluation Settings
For CLIP score (Hessel et al., 2021), we apply the openai-clip  [^1] https://github.com/openai/CLIP library to measure the CLIP score between prompts and generated images. We employ ViT-B/32 as the CLIP model. To measure HPSv2 (Wu et al., 2023a) score, we generate 3.2k images across the four categories (800 images per category) and use the provided human preference predictor to estimate the performance. For the Fréchet Inception Distance (FID) (Heusel et al., 2017) measurement, we apply clean-fid  [^2] https://github.com/GaParmar/clean-fid library to measure the FID between generated images and ground-truth images.
A.5 Hardware and Software Resources
@@
Our code is running under Python 3 using Anaconda virtual environments and open-source repository forks based on Q-Diffusion (Li et al., 2023). We modified the code to implement our FP4DiT and enable the interface between Quantization scripts and the Hugging Face Diffusers  [^3] https://huggingface.co/docs/diffusers/en/index library. The hardware cost measurement is conducted using the ptflops  [^4] https://github.com/sovrasov/flops-counter.pytorch library. We provide a code implementation with README listing all necessary details and steps.
A.6 Additional Visualization Results